OpenStudy (anonymous):

Solve ^3sqrt(2x+5) --4=0

OpenStudy (anonymous):

\[\sqrt[3]{2x+5}-4=0\]\[\sqrt[3]{2x+5}=4\] you can elevate both sides to the third power \[2x+5=4^3\]\[x= {4^3-5 \over 2}\]